OpenStudy (anonymous):

determine the slope of the following curve at the given point. 6xy^(1/2)+y=13; (2,1)

OpenStudy (anonymous):

major help needed

OpenStudy (anonymous):

\[6x\sqrt{y}+y=13\] right?

OpenStudy (anonymous):

written a diff way but yes

OpenStudy (anonymous):

so your job it to find y' using implicit differentiation and evaluate at (2,1)

OpenStudy (anonymous):

the derivative of y is no problem, it is y' and the derivative of 13 is 0 the work is in finding the derivative of \[6x\sqrt{y}\] which requires the product rule

OpenStudy (anonymous):

slope is def not 1

OpenStudy (anonymous):

-6/7

OpenStudy (anonymous):

so taking the derivative wrt x you get \[6\sqrt{y}+6x\times \frac{1}{2\sqrt{y}}\times y'\]

OpenStudy (anonymous):

that is the only complicated step in this problem knowing to use the produce rule and the chain rule. this gives \[6\sqrt{y}+\frac{3y'}{\sqrt{y}}+y'=0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i made a typo hold on

OpenStudy (anonymous):

\[6\sqrt{y}+\frac{3xy'}{\sqrt{y}}+y'=0\]

OpenStudy (anonymous):

i forgot the x. now you can either use algebra now to solve for y' and then replace x by 2 and y by 1 or else you can replace x by 2, y by 1 and then solve for y' either way

OpenStudy (anonymous):

i just need the slope though

OpenStudy (anonymous):

yeah, is it clear what you have to do?

OpenStudy (anonymous):

you want y' when x = 2 and y = 1

OpenStudy (anonymous):

do you understand the whole equation on a tangent line that concept..i do not understand..ill give you an example

OpenStudy (anonymous):

\[6\sqrt{y}+\frac{3xy'}{\sqrt{y}}+y'=0\] so if x = 2, y = 1 you get \[6\sqrt{1}+\frac{3\times 2y'}{\sqrt{1}}+y'=0\] or just \[6+6y'+y'=0\] and solve this for y'

OpenStudy (anonymous):

you get \[y'=-\frac{6}{7}\]

OpenStudy (anonymous):

like determine 2x^2+3xy+3y^2=17;(2,1)

OpenStudy (anonymous):

i got that =)

OpenStudy (anonymous):

right, use implicit diff, then you will get y' in terms of x and y, then replace x and y by the numbers given and you will get the slope

OpenStudy (anonymous):

i always seem to make stupid mistakes

OpenStudy (anonymous):

so in your example step one would be \[4x+3y+3xy'+6yy'=0\] and then either solve for y' directly and substitute, or substitute and solve for y'

OpenStudy (anonymous):

ok..ill try again =(

OpenStudy (anonymous):

its just frustrating lol

OpenStudy (anonymous):

important part is to remember product and chain rule, and also that you are thinking \[y=f(x), y'=f'(x)\] so for example if you see \[\sqrt{x}y\] you think \[\sqrt{x}f(x)\] and if you see \[xy^2\] you think \[xf^2(x)\] that will remind you to use product and chain rule

OpenStudy (anonymous):

did you get -4/3x - 11/12?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

its wrong =(

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

??